# DSDE: Dynamic Speculative Decoding with KLD Stability for Real-World Serving

### v1

@@ Abstract
Index Terms:
Speculative Decoding, LLM Inference, Dynamic Speculation Length, Large-Batch Serving, Kullback–Leibler Divergence
I Introduction
Large Language Models (LLMs) have become central to modern NLP, but their practical deployment in latency-sensitive services remains constrained by slow auto-regressive decoding. Although GPUs provide high throughput, generation proceeds token-by-token, which limits hardware utilization and increases per-sequence latency. LLM inference is fundamentally memory-bound, as each decoding step repeatedly accesses the model weights and the KV cache. This leads to low arithmetic intensity and under-utilization of compute resources [1]. To mitigate this, batching is widely adopted [2, 3], which improves GPU utilization by shifting the bottleneck toward compute. However, batching alone does not reduce the inherent latency of any single sequence and can even introduce queuing delays in latency-sensitive environments [4]. This motivates complementary algorithmic strategies that directly reduce decoding latency. Speculative decoding has emerged as one of the most effective such strategies [5]. A lightweight draft model proposes multiple candidate tokens, and the target model verifies them in parallel. This approach reduces the number of target model inferences and improves decoding efficiency. Higher acceptance rates yield fewer decoding steps and greater speedups. However, current implementations typically use a static speculation length (SL) across all sequences and decoding steps. This rigidity is problematic in heterogeneous large-batch serving: mixed tasks such as code generation and dialogue require different speculation strategies (Table I). A single static SL forces the entire batch to follow the least efficient sequence, creating a scalability bottleneck. This limitation motivates a more flexible approach, in which speculation length is assigned per sequence within a batch.
TABLE I: Performance of Static SL Strategies on Heterogeneous Tasks (HumanEval vs. ShareGPT), showing latency(second) and block efficiency differences between $SL=2$ and $SL=8$.
@@ II Related Work
II-A Architectural Paradigms of Speculative Decoding
@@
II-B Algorithmic Enhancements for Dynamic SL
@@
II-C Stability as a Signal in Language Generation
@@
II-D System-Level Optimizations
@@ III Methodology
Figure 4: The workflow of our dynamic Speculative length (SL) system. After the standard speculative decoding steps involving the Draft worker (1) , Target worker (2) , and Rejection sampler (3) , our SL adapter (4) computes the next SL, which informs the Lookahead scheduler for the subsequent decoding round.
III-A Dynamic SL adapter
III-A1 Dynamic Calibration of Maximum Speculation Length
@@
III-A2 Dynamic Speculation Length Decoding Algorithm
@@
III-B Per-sequence speculative decoding in vLLM
@@
III-C Adaptive Speculative Length Capping
@@
IV Experiments
IV-A Experimental Setup
@@
IV-B Evaluation of Predictive Signals
@@
Table II presents the Pearson correlation coefficients between these signals and the token acceptance rate at sampling temperatures of 0.0 and 1.0. The results show that all evaluated signals exhibit a weak correlation with token-level acceptance, and this relationship weakens further when moving from near-greedy decoding to stochastic sampling.
TABLE II: Pearson correlation (r) between signals and token acceptance probability at different sampling temperatures on the CNN/DM dataset.
@@ IV-C End-to-End Performance and Robustness
Table III reports results for the LLaMA-3.1-70B-Instruct (target) with LLaMA-3.2-1B-Instruct (draft). At temperature 0.0, all accelerated methods achieved substantial speedups over the incremental baseline (38.41s). Our algorithm reached 13.97s, comparable to AdaEDL (13.83s) and close to static-opt (13.44s), but without profiling overhead.
@@
TABLE III: Latency results and speed-ups compared to the autoregressive decoding for LLaMA3.1-70B-Instruct (target) with LLaMA3.2-1B-Instruct (draft). Values are in seconds.
Method Mean Latency (s) Speedups
Autoregressive 38.41 1.00x
Static-opt 13.44 2.86x
Proposed Dynamic SL 13.97 2.75x
AdaEDL (base=7) 13.83 2.78x
Method Mean Latency (s) Speedups
Autoregressive 38.47 1.00x
Static-opt 18.02 2.13x
Proposed Dynamic SL 19.19 2.00x
AdaEDL (base=7) 17.64 2.17x
@@
Figure 7: Mean latency of the WVIR-based algorithm compared to AdaEDL and the per-dataset Static-opt baseline across dataset at temperature $0.$
IV-D Analysis in a Low-Acceptance-Rate Regime
@@
The relative performance degradation is detailed in Table IV, which presents the latency for the Gemma model pair, normalized by the latency achieved with the LLaMa3.1-70B/LLaMa3.2-1B model pair. The results reveal a notable divergence in performance under these adverse conditions. While the proposed WVIR-based algorithm maintained a latency close to the Static-opt baseline (e.g., 234% the static-opt latency on CNNDM), AdaEDL’s performance degraded substantially.
@@
TABLE IV: Percentile increment in mean latency of dynamic methods with Gemma 27B/2B, compared to the Llama-70B/1B.
@@
IV-E Efficacy of Per-Sequence Adaptation and $SL_{cap}$
This section evaluates the scalability of the per-sequence approach and the efficacy of the $SL_{cap}$ in mitigating its limitations. To conduct this experiment, we used the CNN/DM dataset and tested batch sizes from 1 to 64, measuring the corresponding throughput in tokens per second. In an ideal scalable system, throughput should increase linearly with batch size. However, the results in Figure 9 demonstrate that the naive per-sequence strategy (Dynamic SL (No Cap)) exhibits suboptimal scalability as the batch size increases. This scalability gap is consistent across temperatures. Compared to its performance at a batch size of 1, the No Cap strategy’s throughput scales by only 11.21x at temperature=0.0 and 11.92x at temperature=1.0. This degradation is caused by the ”straggler problem” as mentioned in Section I: a few sequences with aggressively long predicted speculation lengths stall the entire batch. This finding highlights that while a naive per-sequence approach is theoretically appealing, it suffers from practical scalability issues in large-batch environments.

### v3

@@ Abstract
22footnotetext: These authors contributed equally.11footnotetext: Corresponding author: ej85.jeon@samsung.com33footnotetext: This paper has been accepted for presentation at the 11th Special Session on Intelligent Data Mining, IEEE International Conference on Big Data (BigData 2025).44footnotetext: © 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting or republishing for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
Keywords: Speculative Decoding, LLM Inference, Dynamic Speculation Length, Large-Batch Serving, Kullback–Leibler Divergence
1 Introduction
Large Language Models (LLMs) have become central to modern NLP, but their practical deployment in latency-sensitive services remains constrained by slow auto-regressive decoding. Although GPUs provide high throughput, generation proceeds token-by-token, which limits hardware utilization and increases per-sequence latency. LLM inference is fundamentally memory-bound, as each decoding step repeatedly accesses the model weights and the KV cache. This leads to low arithmetic intensity and under-utilization of compute resources [1]. To mitigate this, batching is widely adopted [2, 3], which improves GPU utilization by shifting the bottleneck toward compute. However, batching alone does not reduce the inherent latency of any single sequence and can even introduce queuing delays in latency-sensitive environments [4]. This motivates complementary algorithmic strategies that directly reduce decoding latency. Speculative decoding has emerged as one of the most effective such strategies [5]. A lightweight draft model proposes multiple candidate tokens, and the target model verifies them in parallel. This approach reduces the number of target model inferences and improves decoding efficiency. Higher acceptance rates yield fewer decoding steps and greater speedups. However, current implementations typically use a static speculation length (SL) across all sequences and decoding steps. This rigidity is problematic in heterogeneous large-batch serving: mixed tasks such as code generation and dialogue require different speculation strategies (Table 1). A single static SL forces the entire batch to follow the least efficient sequence, creating a scalability bottleneck. This limitation motivates a more flexible approach, in which speculation length is assigned per sequence within a batch.
Table 1: Performance of Static SL Strategies on Heterogeneous Tasks (HumanEval vs. ShareGPT), showing latency (second) and block efficiency (BE) differences between $SL=2$ and $SL=8$.
@@ 2 Related Work
2.1 Architectural Paradigms of Speculative Decoding
@@
2.2 Algorithmic Enhancements for Dynamic SL
@@
2.3 Stability as a Signal in Language Generation
@@
2.4 System-Level Optimizations
@@ 3 Methodology
Figure 4: Workflow of the dynamic Speculative Length (SL) system. After the standard speculative decoding steps involving the Draft worker (1), Target worker (2), and Rejection sampler (3), the SL adapter (4) computes the next SL, which informs the Lookahead scheduler for the subsequent decoding round.
3.1 Dynamic SL adapter
3.1.1 Dynamic Calibration of Maximum Speculation Length
@@
3.1.2 Dynamic Speculation Length Decoding Algorithm
@@
3.2 Per-sequence speculative decoding in vLLM
@@
3.3 Adaptive Speculative Length Capping
@@
4 Experiments
4.1 Experimental Setup
@@
4.2 Evaluation of Predictive Signals
@@
Table 2 presents the Pearson correlation coefficients between these signals and the token acceptance rate at sampling temperatures of 0.0 and 1.0. The results show that all evaluated signals exhibit a weak correlation with token-level acceptance, and this relationship weakens further when moving from near-greedy decoding to stochastic sampling.
Table 2: Pearson correlation ($r$) between signals and token acceptance probability at different sampling temperatures on the CNN/DM dataset.
@@ 4.3 End-to-End Performance and Robustness
Table 3 reports results for the LLaMA-3.1-70B-Instruct (target) with LLaMA-3.2-1B-Instruct (draft). At temperature 0.0, all accelerated methods achieved substantial speedups over the incremental baseline (38.41s). Our algorithm reached 13.97s, comparable to AdaEDL (13.83s) and close to static-opt (13.44s), but without profiling overhead.
@@
Table 3: Latency results and speed-ups compared to the autoregressive decoding for LLaMA3.1-70B-Instruct (target) with LLaMA3.2-1B-Instruct (draft). Values are in seconds.
(a) Temperature 0.0
Method Mean Latency (s) Speedup Autoregressive 38.41 1.00× Static-opt 13.44 2.86× Proposed Dynamic SL 13.97 2.75× AdaEDL (base=7) 13.83 2.78×
(b) Temperature 1.0
Method Mean Latency (s) Speedup Autoregressive 38.47 1.00× Static-opt 18.02 2.13× Proposed Dynamic SL 19.19 2.00× AdaEDL (base=7) 17.64 2.17×
@@ 4.4 Analysis in a Low-Acceptance-Rate Regime
The relative performance degradation is detailed in Table 4, which presents the latency for the Gemma model pair, normalized by the latency achieved with the LLaMa3.1-70B/LLaMa3.2-1B model pair. The results reveal a notable divergence in performance under these adverse conditions. While the proposed WVIR-based algorithm maintained a latency close to the Static-opt baseline (e.g., 234% the static-opt latency on CNNDM), AdaEDL’s performance degraded substantially.
Figure 7: Mean latency of the WVIR-based algorithm compared to AdaEDL and the per-dataset Static-opt baseline across datasets at temperature $0.$
@@
Table 4: Percentile increment in mean latency of dynamic methods with Gemma 27B/2B, compared to the LLaMA-70B/1B.
@@
4.5 Efficacy of Per-Sequence Adaptation and $SL_{cap}$
This section evaluates the scalability of the per-sequence approach and the efficacy of the $SL_{cap}$ in mitigating its limitations. To conduct this experiment, we used the CNN/DM dataset and tested batch sizes from 1 to 64, measuring the corresponding throughput in tokens per second. In an ideal scalable system, throughput should increase linearly with batch size. However, the results in Figure 9 demonstrate that the naive per-sequence strategy (Dynamic SL (No Cap)) exhibits suboptimal scalability as the batch size increases. This scalability gap is consistent across temperatures. Compared to its performance at a batch size of 1, the No Cap strategy’s throughput scales by only 11.21x at temperature=0.0 and 11.92x at temperature=1.0. This degradation is caused by the ”straggler problem” as mentioned in Section 1: a few sequences with aggressively long predicted speculation lengths stall the entire batch. This finding highlights that while a naive per-sequence approach is theoretically appealing, it suffers from practical scalability issues in large-batch environments.